OpenStudy (anonymous):

Choose the equation below that represents the line passing through the point (-2, -3) with a slope of -6

OpenStudy (anonymous):

A) y + 3 = -6(x + 2) B) y + 2 = 6(x + 3) C) y - 3 = -6(x - 2) D) y - 2 = 6(x - 3)

OpenStudy (phi):

if you had an equation in the form y = m*x + b m (the number multiplying x) is the slope so if you change the first equation by multiplying the -6 times each term in the parens y + 3 = -6(x + 2) y+3 = -6x-12 and add -3 to both sides y+3-3= -6x-12-3 which simplifies to y= -6x-15 so option A 's equation has a slope of -6. It is a possible answer to see if (-2,-3) works, replace x with -2, and see if you get y= -3 can you do that?

OpenStudy (anonymous):

no just give me the answer